OpenStudy (anonymous):

Using the time constant at noon, with the temperature in professor snaf's office at 75(f) and the ouside temp at 95(f), the ac breaks down. Suppose the time constant for professor snarf's office is 4 hours. A) determine the temp in professors snarf's office at 2:oo pm B) Determine when the temperature in professor snarf's office will reach 80 (f)

OpenStudy (anonymous):

Still looking for help with this. I have two equations I've been able to do nothing with. The values are not working out for the equations. \[\frac{DT}{dt}=-k(T-M)\] and \[T(t)=T_o e^{-kt}+M(1-e^{-kt})\]